OpenStudy (anonymous):

Mona can fill a tank in 4 hours,if Lisa helps her it only takes them 2.5 hrs. how long will it take lisa to fill the tank alone?

OpenStudy (anonymous):

in 1 hr mona can fill 1/4 of the tank where as they together in one 1 hr can fill 1/2.5 =2/5 of the tank

OpenStudy (anonymous):

thus in 1hr Lisa alone can fill (2/5 -1/4)=3/20 part of the tank hence lisa can fill the tank alone in 20/3 hrs

OpenStudy (anonymous):

thank you...